# Birds look like cars: Adversarial analysis of intrinsically interpretable deep learning

### v1

@@ 5 Discussion
Conceptually, CBMs look more interpretable and robust than prototype-based networks. First, they remove the prototype projection stage in favor of labeled concepts that cannot be manipulated in our scenario considering ProtoViT. Intuitively, restricting the set of concepts entails less opportunity for shortcut learning (or at least the possible shortcuts will be named) and more opportunity for human-guided model correction (Shin et al., 2023; Kim et al., 2024). Furthermore, CBMs require no self-supervised pre-training of prototypes to align similarity between latent representations with the human visual perception of similarity like in PIP-Net. CBMs have a predefined, controllable sparsity that can potentially circumvent the representation redundancy issue in prototype-based networks (Turbé et al., 2024).
Future work on adversarial analysis of intrinsically interpretable deep learning should explore the vulnerabilities of CBMs. There currently is no work on backdoor attacks against CBMs, and whether these could be disguised. In general, worth exploring are semantic triggers like rotation, change of color temperature, or sharpness, as well as defense mechanisms against adversarial manipulation (see Noppel and Wressnegger, 2024, and references given there). ProtoPDebug (Bontempelli et al., 2023) could potentially be leveraged as a human-level defense mechanism against adversarial manipulation.
@@
Recent work in this direction explores several evaluation metrics for ProtoViTs (Turbé et al., 2024) like completeness, contrastivity, consistency, and stability, which may correlate with the networks’ robustness. In a concurrent work, Pach et al. (2025) propose aligning prototypical parts with object parts and separating color and non-color visual features, which improves the interpretability of prototype-based convolutional networks. While our paper presents an empirical case, future work could deepen the theoretical analysis of why interpretable networks are vulnerable and how the attacks succeed (akin Dombrowski et al., 2019; Leemann et al., 2023; Baniecki et al., 2024).
@@ A.1 Datasets and preprocessing
CUB-200-201 [Wah et al., 2011] contains 11788 images of 200 bird species (classes), split into 5994 for training and 5794 for testing. We follow prior work on prototype-based networks [Chen et al., 2019, Ma et al., 2024] to apply standard data augmentation (rotate, skew, shear, flip) and preprocessing steps (normalization), including cropping the images to bounding boxes and resizing them to $224\times 224$. Stanford Cars [Krause et al., 2013] contains 16185 images of 196 car classes, where we only use the training set with 8144 images, which are resized to $224\times 224$.
Out-of-distribution Birds is our custom-made set of 2044 images from 13 bird species (classes) coming from the training set of ImageNet [Russakovsky et al., 2015]. Crucially, some ImageNet classes overlap with CUB (e.g. 13 – junco, 16 – bulbul, 94 – hummingbird, 144 – pelican, 146 – albatross), and thus we collected 200 images from each of the following classes that we did not find in CUB: 7 – cock, 8 – hen, 17 – jay, 18 – magpie, 19 – chickadee, 21 – kite, 22 – bald eagle, 23 – vulture, 24 – great grey owl, 127 – white stork, 128 – black stork, 134 – crane, 145 – king penguin. We then manually filtered the resulting 2600 images to remove images of low quality, with objects (birds) occupying a very small portion of an image, or labeling errors (e.g. Fig. 9).
@@
ISIC 2019 [Tschandl et al., 2018, Hernández-Pérez et al., 2024] contains 33569 dermoscopic images of 9 classes, split into 25331 for training and 8238 for testing. Since prior work [Nauta et al., 2023a] provides no code to reproduce the results, we rely on the data augmentation and preprocessing pipeline from PIP-Net [Nauta et al., 2023b], including using TrivialAugmentWide [Müller and Hutter, 2021] and resizing images to $224\times 224$. To simplify the analysis, we treat the ‘melanocytic nevus (NV)’ class as benign and combine the remaining 8 abnormal (e.g. pre-cancerous, cancerous) classes as ‘malignant’. Thus, the training set has a 50% ratio in binary outcomes; the test set has 70% due to 25% images of class ‘unknown’ being included in ‘malignant’ so as to simulate the need for human oversight. MURA [Rajpurkar et al., 2018] contains 40005 musculoskeletal X-ray images of 2 classes (either normal or not), split into 36808 for training and 3197 for testing. Data preprocessing is similar to ISIC, taking into account the grayscale nature of images (refer to supplementary code for details). The training set has a 40% ratio in binary outcomes, and the test set has 48%.
@@ A.2 Training
We rely on the default hyperparameters available in code implementation for training ProtoViT. We set $\tau=100$, $\lambda_{\mathrm{cross-entropy}}=1.0$, $\lambda_{\mathrm{cluster}}=-0.8$, $\lambda_{\mathrm{separation}}=0.1$, $\lambda_{L_{1}}=1e{-}2$, $\lambda_{\mathrm{orthogonality}}=1e{-}3$, $\lambda_{\mathrm{coherence}}=3e{-}3$ and $\lambda_{\mathrm{coherence}}=1e{-}6$ for slots pruning. The warm-up stage takes 5 epochs with learning rates $\mathrm{lr}_{f}=1e{-}7$ and $\mathrm{lr}_{g}=3e{-}3$, joint stage takes 10 epochs with $\mathrm{lr}_{f}=5e{-}5$ and $\mathrm{lr}_{g}=3e{-}3$, slots pruning takes 5 epochs with $\mathrm{lr}_{g}=5e{-}5$, fine-tuning takes 10 epochs with $\mathrm{lr}_{h}=1e{-}4$. We use the AdamW optimizer and a batch size of $128$.
@@
We rely on the default hyperparameters available in code implementation for training PIP-Net. For training, we set $\lambda_{\mathrm{C}}=2$, $\lambda_{\mathrm{A}}=5$, $\lambda_{\mathrm{U}}=2$. We modify these parameters accordingly for backdoor fine-tuning (see Section 4.2). The pre-training stage takes 5 epochs when the classification layer $h$ is frozen, then 3 epochs consider fine-tuning only layer $h$, then 7 epochs consider freezing only the encoder backbone (part of $f$), and the next 25 epochs of training the whole model. Learning rates are set to $\mathrm{lr}_{f}=5e{-}4$ and $\mathrm{lr}_{h}=5e{-}2$ in a cosine annealing schedule. We use the AdamW optimizer and a batch size of $64$ due to forwarding $2\times 64$ inputs through the network at each step to calculate the alignment loss.

### v2

@@ 6 Discussion
Conceptually, CBMs look more interpretable and robust than prototype-based networks. First, they remove the prototype projection stage in favor of labeled concepts that cannot be manipulated in our scenario, considering ProtoViT. Intuitively, restricting the set of concepts entails less opportunity for shortcut learning (or at least the possible shortcuts will be named) and more opportunity for human-guided model correction (Shin et al., 2023; Kim et al., 2024). Furthermore, CBMs require no self-supervised pre-training of prototypes to align the similarity between latent representations with the human visual perception of similarity, like in PIP-Net. CBMs have a predefined, controllable sparsity that can potentially circumvent the representation redundancy issue in prototype-based networks (Turbé et al., 2024).
Future work on adversarial analysis of intrinsically interpretable deep learning should explore the vulnerabilities of CBMs. There is currently no work on backdoor attacks against CBMs, and whether these could be disguised. In general, worth exploring are semantic triggers like rotation, change of color temperature, or sharpness, as well as defense mechanisms against adversarial manipulation (see Noppel and Wressnegger, 2024, and references given there). ProtoPDebug (Bontempelli et al., 2023) could potentially be leveraged as a human-level defense mechanism against adversarial manipulation.
@@
Recent work in the direction of aligned interpretable models explores several evaluation metrics for ProtoViTs (Turbé et al., 2024) like completeness, contrastivity, consistency, and stability, which may correlate with the networks’ robustness. In a concurrent work, Pach et al. (2025) propose aligning prototypical parts with object parts and separating color and non-color visual features, which improves the interpretability of prototype-based convolutional networks. While our paper presents an empirical case, future work could deepen the theoretical analysis of why interpretable networks are vulnerable and how the attacks succeed (akin Dombrowski et al., 2019; Leemann et al., 2023; Baniecki et al., 2024).
@@ A.1 Datasets and preprocessing
CUB-200-2011 [Wah et al., 2011] contains 11788 images of 200 bird species (classes), split into 5994 for training and 5794 for testing. Each class has about 30 samples in the training and test sets. We follow prior work on prototype-based networks [Chen et al., 2019, Ma et al., 2024] to apply standard data augmentation (rotate, skew, shear, flip) and preprocessing steps (normalization), including cropping the images to bounding boxes and resizing them to $224\times 224$. Stanford Cars [Krause et al., 2013] contains 16185 images of 196 car classes, where we only use the training set with 8144 images, which are resized to $224\times 224$.
Out-of-distribution Birds is our custom-made set of 2044 images from 13 bird species (classes) coming from the training set of ImageNet [Russakovsky et al., 2015]. Crucially, some ImageNet classes overlap with CUB, e.g. 13 – junco, 16 – bulbul, 94 – hummingbird, 144 – pelican, 146 – albatross, and thus we collected 200 images from each of the following classes that we did not find in CUB: 7 – cock, 8 – hen, 17 – jay, 18 – magpie, 19 – chickadee, 21 – kite, 22 – bald eagle, 23 – vulture, 24 – great grey owl, 127 – white stork, 128 – black stork, 134 – crane, 145 – king penguin. We then manually filtered the resulting 2600 images to remove images of low quality, with objects (birds) occupying a very small portion of an image, or labeling errors, e.g. Figure 9.
@@
ISIC 2019 [Tschandl et al., 2018, Hernández-Pérez et al., 2024] contains 33569 dermoscopic images of 9 classes, split into 25331 for training and 8238 for testing. Since prior work [Nauta et al., 2023a] provides no code to reproduce the results, we rely on the data augmentation and preprocessing pipeline from PIP-Net [Nauta et al., 2023b], including using TrivialAugmentWide [Müller and Hutter, 2021] and resizing images to $224\times 224$. To simplify the analysis, we treat the ‘melanocytic nevus (NV)’ class as benign and combine the remaining 8 abnormal, e.g. pre-cancerous, cancerous, classes as ‘malignant’. Thus, the training set has a 50% ratio in binary outcomes; the test set has 70% due to 25% images of class ‘unknown’ being included in ‘malignant’ so as to simulate the need for human oversight. MURA [Rajpurkar et al., 2018] contains 40005 musculoskeletal X-ray images of 2 classes (either normal or not), split into 36808 for training and 3197 for testing. Data preprocessing is similar to ISIC, taking into account the grayscale nature of images (refer to supplementary code for details). The training set has a 40% ratio in binary outcomes, and the test set has 48%.
@@ A.2 Training
We rely on the default hyperparameters available in the code implementation for training ProtoViT. We set $\tau=100$, $\lambda_{\mathrm{cross-entropy}}=1.0$, $\lambda_{\mathrm{cluster}}=-0.8$, $\lambda_{\mathrm{separation}}=0.1$, $\lambda_{L_{1}}=1e{-}2$, $\lambda_{\mathrm{orthogonality}}=1e{-}3$, $\lambda_{\mathrm{coherence}}=3e{-}3$ and $\lambda_{\mathrm{coherence}}=1e{-}6$ for slots pruning. The warm-up stage takes 5 epochs with learning rates $\mathrm{lr}_{f}=1e{-}7$ and $\mathrm{lr}_{g}=3e{-}3$, joint stage takes 10 epochs with $\mathrm{lr}_{f}=5e{-}5$ and $\mathrm{lr}_{g}=3e{-}3$, slots pruning takes 5 epochs with $\mathrm{lr}_{g}=5e{-}5$, fine-tuning takes 10 epochs with $\mathrm{lr}_{h}=1e{-}4$. We use the AdamW optimizer and a batch size of $128$.
@@
We rely on the default hyperparameters available in the code implementation for training PIP-Net. For training, we set $\lambda_{\mathrm{C}}=2$, $\lambda_{\mathrm{A}}=5$, $\lambda_{\mathrm{U}}=2$. We modify these parameters accordingly for backdoor fine-tuning (see Section 5.2). The pre-training stage takes 5 epochs when the classification layer $h$ is frozen, then 3 epochs consider fine-tuning only layer $h$, then 7 epochs consider freezing only the encoder backbone (part of $f$), and the next 25 epochs of training the whole model. Learning rates are set to $\mathrm{lr}_{f}=5e{-}4$ and $\mathrm{lr}_{h}=5e{-}2$ in a cosine annealing schedule. We use the AdamW optimizer and a batch size of $64$ due to forwarding $2\times 64$ inputs through the network at each step to calculate the alignment loss.